# Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks

### 22 May 2020 v1

@@ 3.2 Abstractive Question Answering
Because RAG leverages an encoder-decoder model, it can go beyond extractive question-answering and answer questions with free-form, abstractive text generation. To test RAG’s ability to generate natural language responses in a knowledge-intensive setting, we use the MS-MARCO Natural Language Generation task v2.1 [38]. This task consists of natural language questions submitted to a search engine, ten snippets retrieved from a search engine for each question, and a full sentence natural language answer annotated from these retrieved passages.
As we are interested in models that can perform their own latent retrieval, we do not use the supplied passages, only the questions and answers, thus treating MS-MARCO as an open-domain abstractive question answering task. MS-MARCO does contain some questions that cannot be answered in a way that matches the reference answer without access to the context passages, such as “What is the weather in volcano, CA?” so we note that performance on Open-MSMARCO will be lower than models that do use these gold context passages.
We further note that there are questions in MS-MARCO that cannot be answered using a Wikipedia knowledge source alone. In these cases, RAG can rely on the parametric implicit knowledge in its BART parameters in order to generate commonsense responses.
@@ 3.3 Jeopardy Question Generation
In order to further evaluate RAG’s generation abilities in a non-question answering setting, we propose to study Open-domain question generation. Rather than repurpose questions from standard open-domain QA tasks, which typically consist of short and simple questions, we instead propose to study the more demanding task of generating of Jeopardy questions. Jeopardy is an unusual format that consists of trying to guess an entity from a fact about that entity. For example, “The World Cup” is the answer to the jeopardy question “In 1986 Mexico scored as the first country to host this international sports competition twice.” As Jeopardy “questions" are precise, factual statements, generating Jeopardy-style questions conditioned on the answer entity they refer to constitutes a challenging knowledge-intensive generation task.
We use the raw Jeopardy data and splits from SearchQA [10], consisting of 97,391 training, 13,713 development, and 26,848 test datapoints. As this is a new task, we also train a BART system to compare RAG to. Following [60], we evaluate generations using the SQuAD-tuned Q-BLEU-1 metric [37]. Q-BLEU-1 is a variant of BLEU-1 which puts a higher weight on the matching entities, and has higher correlation with human judgment for question generation compared to standard word overlap metrics.
As automatic metrics can be unreliable, especially on such open-ended tasks, we also perform a human evaluation of generations. We run two evaluations, one to assess the factuality of generations, and one to assess specificity. We follow the recent best-practice of performing a pairwise comparative evaluation between two systems [30]. Assessors are shown an answer entity and two generated questions about that entity, one from BART and one from RAG. They are then asked to pick one of four possible options—Sentence A is better, Sentence B is better, both are correct or neither is good.
@@ 3.4 Fact Verification
FEVER [50] is a fact verification dataset that involves classifying whether a natural language claim is supported or refuted by Wikipedia, or whether there is not enough information to decide. The task requires retrieving evidence from Wikipedia relating to the claim and then reasoning about the retrieved evidence to classify whether the claim is true, false, or unverifiable from Wikipedia alone. FEVER is a retrieval problem coupled with an entailment reasoning task. It also provides a good test bed for exploring the RAG models’ ability to handle classification rather than generation.
We map FEVER class labels (supports, refutes, or not enough info) to single output tokens and directly train with claim-class pairs. Crucially, unlike most other approaches to FEVER, we do not use supervision on retrieved evidence. We explore two different FEVER variants: the standard 3-way classification task (supports/refutes/not enough info) and the 2-way FEVER (supports/refutes) task studied in Thorne and Vlachos [51]. In both cases we report label accuracy.
3.5 Implementation Details
@@ 4 Results
Open-Book REALM [18] 40.4 - / - 40.7 46.8
@@
RAG-Sequence 44.5 56.1 / 68.0 45.2 52.2
Table 1: Open-Domain QA Test Scores. For TQA, the left column uses the test split commonly used in Open-Domain QA. The right column uses the hidden TQA Wiki test split. See Appendix B for further information.
Model Jeopardy QGen MS-MARCO FEVER-3 FEVER-2
B-1 QB-1 R-L B-1 Label Accuracy
@@
RAG-Token 17.3 22.2 40.1 41.5 72.5 89.5
RAG-Sequence 14.7 21.4 40.8 44.2
Table 2: Generation and classification task Test Scores. SotA for MS-MARCO is [4], FEVER-3 is [61] and FEVER-2 is [51] * Uses gold context/evidence, best-performing model without gold access underlined. As FEVER is a classification dataset, RAG-Token and RAG-Sequence are equivalent.
@@ 4.1 Open-domain Question Answering
Table 1 shows results for RAG along with recent state-of-the-art models. On all four open-domain QA tasks, RAG sets a new state-of-the-art (in the case of TQA only on the T5-comparable split).
RAG combines the generation flexibility of the “closed-book" (parametric only) approaches and the performance of "open-book" retrieval-based approaches. Unlike REALM and T5+SSM, RAG enjoys strong results without expensive specialized "salient span masking" pre-training [18], relying on off-the-shelf components. It is worth noting that RAG’s retriever is initialized using DPR’s retriever, which does use retrieval supervision on Natural Questions and TriviaQA. RAG compares favourably to DPR QA system on open-domain QA, which uses a BERT-based cross-encoder system to re-rank documents, along with an extractive reader. RAG demonstrates that neither a re-ranker nor extractive reader is necessary for state-of-the-art machine reading performance. Generating answers even when it is possible to extract them has a number of advantages. Documents which contain clues as to the correct answer but do not contain the correct answer verbatim themselves can still contribute probability mass towards a correct answer being generated, which is not possible with standard extractive approaches, leading to more effective marginalization across documents. Furthermore, RAG is able to generate correct answers even when the correct answer is not present in any of the retrieved documents, achieving an accuracy of 11.8% in such cases for Natural Questions, whereas an extractive model would score 0%.
@@ 4.2 Abstractive Question Answering
As shown in Table 2, RAG-Sequence outperforms BART on Open MS-MARCO generation by 2.6 Bleu points and 2.6 Rouge-L points. It approaches the performance of state-of-the-art models, which is impressive considering that (i) these models have access to passages that contain the specific information required to generate the reference answer, (ii) many questions are unanswerable without access to gold passages, and (iii) other questions are unanswerable from Wikipedia alone. Table 4 shows some generated answers from our models. Qualitatively, we find that RAG models hallucinate less and generate factually correct text more often than BART. Later we also show that RAG generations are more diverse than BART generations (see Section 4.6).
@@ 4.3 Jeopardy Question Generation
Table 2 shows automatic metric results on the Jeopardy question generation task. We find that RAG-Token performs better than the RAG-Sequence model in this setting, with both models outperforming BART using the Q-BLEU-1 metric.
Table 3 shows the results from the human evaluation. The human evaluation was carried out with 452 pairs of generations from BART and RAG-Token. The annotators indicated that BART was more factual than RAG in only 7.1% of cases, while RAG was more factual in 42.7% of cases and both RAG and BART were factual in a further 17% of cases, clearly demonstrating the comparative effectiveness of RAG on the task over a state-of-the-art conditional generation model. The annotators also strongly prefer RAG generations in terms of specificity.
BART better RAG-Token better Both good Both poor No Majority
Factuality 7.1% 42.7% 11.7% 17.7% 20.8%
Specificity 16.8% 37.4% 18.8% 6.9% 20.1%
Typical example of generations from each model are shown in Table 4. BART generates a more generic response (which is incorrect), whereas the RAG models generate specific and correct facts about Washington state.
We hypothesise that RAG-Token performs best for this task as Jeopardy questions often contain two separate pieces of information about the entity, and RAG-Token is able to synthesize a response by combining disparate information from different retrieved documents in one generation. Figure 2 shows an example where content from two documents has been combined to produce the generated question. Document 2 contains information about Hemingway’s “The Sun also rises,” and the contribution for “Sun” is very high for document 2. Similarly, “A Farewell to Arms” is mentioned in Document 1, which dominates the posterior when this title is generated. Intriguingly, after the first token of these book titles are generated, the distribution over documents flattens again. This observation suggests that the generator completes the book titles without depending on specific documents. In other words, the model’s parametric knowledge is sufficient to complete the titles.
We show evidence for the above interpretation by feeding the BART-only baseline with the partial decoding "The Sun. BART completes the generation "The Sun Also Rises" is a novel by this author of "The Sun Also Rises" indicating the title "The Sun Also Rises" is stored in BART’s parameters. Similarly, feeding the partial decoding "The Sun Also Rises" is a novel by this author of "A will result in BART completing the generation with "The Sun Also Rises" is a novel by this author of "A Farewell to Arms. This example shows how the parametric and non-parametric memories work together—the non-parametric component helps to guide the generation in a particular direction, drawing out specific knowledge stored in the parametric memory.
Figure 2: RAG-Token document posterior $p(z_{i}|x,y_{i},y_{-i})$ for each generated token for input “Hemingway" for Jeopardy generation with 5 retrieved documents. The posterior for document 1 is high when generating “A Farewell to Arms" and for document 2 when generating “The Sun Also Rises"
@@
Table 4: Example Generations for MS-MARCO and Jeopardy Question generation. RAG models generate mpre specific and factually accurate responses, whereas BART generate more factually incorrect (marked by ‘?’), or partially correct (marked by *) and more generic responses.
@@ 4.4 Fact Verification
Table 2 shows our results on the FEVER 3-way and 2-way classification task. For 3-way classification, RAG achieves accuracies that are within 4.3% of state-of-the-art models, which are complex pipeline systems with domain-specific architectures and substantial engineering, trained using intermediate supervision, which RAG does not require.
For 2-way classification, we compare against the model from [51], which trains RoBERTa  [31] to classify the claim as true or false given the gold evidence sentence. RAG achieves an accuracy within 2.7% of this model, despite being supplied with only the claim and retrieving its own evidence.
We also analyze whether the documents retrieved by RAG correspond to the documents annotated as gold evidence in FEVER. We analyze the overlap in Wikipedia articles between the top $k$ documents retrieved by RAG and the gold, annotated evidence documents. We find that the top article retrieved by RAG is a gold document for the claim in 71% of cases, and a gold article is present in the top 10 retrieved articles in 90% of cases.
4.5 Ablations
@@
RAG-Seq 44.0 55.8 44.9 53.4 15.3 21.5 57.2 47.5
Table 5: Ablations on the development set. As FEVER is a classification dataset, RAG-Token and RAG-Sequence are equivalent.
To gain a better understanding of what factors affect RAG’s performance, we perform a number of ablation experiments for our tasks on their respective development sets.
Using more documents
Models are trained with either 5 or 10 retrieved latent documents, and we do not observe significant differences in performance between them. We also have the flexibility to adjust the number of retrieved documents at test time, which does affect performance. Figure 3 (left) shows that retrieving more documents at test time monotonically improves Open-domain QA results for RAG-Sequence, but performance peaks for RAG-Token at 10 retrieved documents. Figure 3 (right) shows that retrieving more documents leads to higher Rouge-L for RAG-Token at the expense of Bleu-1, but the effect is less pronounced for RAG-Sequence.
Figure 3: Left: NQ performance as more documents are retrieved. Center: Fraction of answers in NQ where the answer occurs somewhere in the top K documents. Right: MS-MARCO Bleu-1 and Rouge-L as more documents are retrieved.
Retrieval
@@
A key feature of RAG is the ability to learn to retrieve relevant information for the task at hand. To assess the effectiveness of the retrieval mechanism, we run ablations on RAG where we prevent gradients from propagating into the retriever. Table 5 shows the results across all tasks. In each case, learned retrieval improves results, with the largest improvements in question answering. Figure 3 (center) shows that the learned retriever shows a higher recall for gold documents compared to the fixed retriever. The improvements on TriviaQA and Natural Questions are notable, as we initialize the retriever from DPR, which is trained with strong, document-level supervision to perform well on these tasks. We also compare RAG’s dense embedding-based retrieval mechanism to a word overlap-based BM25 retriever [47]. Here, we replace RAG’s differentiable retriever with a fixed BM25 system. We use the BM25 retrieval scores as logits when calculating $p_{\eta}(z_{i}|x)$. Table 5 and Figure 3 show the results. For FEVER, we find that BM25 performs best, perhaps since FEVER claims are heavily entity-centric and thus well-suited for word overlap-based retrieval. On all other tasks, we find the differentiable retrieval to be helpful, especially question answering, where it is crucial.
4.6 Generation Diversity
Dataset Gold BART RAG-Token RAG-Sequence
MSMARCO 89.6% 70.7% 77.8% 83.5%
Jeopardy Generation 90.0% 32.4% 46.8 % 53.8%
Table 6: Ratio of distinct tri-grams to total tri-grams in the development set generations for MSMARCO and Jeopardy Question Generation.
Section 4.3 established that RAG models generate are more factual and specific than BART for Jeopardy question generation. Similar to Li et al. [29], Vijayakumar et al. [53] and Massarelli et al. [35], we also investigate the diversity of generations by calculating the ratio of distinct ngrams to total ngrams generated by different models. Table 6 shows that RAG-Sequence generations are more diverse than RAG-Token generations, and both generate significantly more diverse outputs than BART without requiring any diversity-promoting decoding strategy.
4.7 Hot-swapping indices
@@
An advantage of non-parametric knowledge models such as RAG is that the knowledge base can be easily updated at test time. Parametric-only models such as T5 or BART require additional training to update their behavior as facts about the world change. As a demonstration, we build an index using the DrQA Wikipedia dump [5], (dated December 21st, 2016) and compare generations from RAG using this index to the newer index used in our main results (December 20th, 2018). We prepared a list of 82 heads of states who had changed between these dates and used a template “Who is {position}?” (e.g., “Who is the prime minister of the UK?”) to query our Natural Questions -finetuned RAG model with each index. RAG achieved an accuracy of 70% using the 2016 index for 2016 world leaders and an accuracy of 68% using the 2018 index for the 2018 world leaders. Only 21% of the model’s predictions were the same using the two indices, and accuracy using mismatched indices is very low (12% using the 2018 index for 2016 leaders and 4% using the 2016 index for 2018 leaders). Our result shows that we can effectively update RAG’s behavior with new world knowledge by simply replacing its non-parametric memory.

### 07 Dec 2020 v2

@@ 3.2 Abstractive Question Answering
RAG models can go beyond simple extractive QA and answer questions with free-form, abstractive text generation. To test RAG’s natural language generation (NLG) in a knowledge-intensive setting, we use the MSMARCO NLG task v2.1 [43]. The task consists of questions, ten gold passages retrieved from a search engine for each question, and a full sentence answer annotated from the retrieved passages. We do not use the supplied passages, only the questions and answers, to treat MSMARCO as an open-domain abstractive QA task. MSMARCO has some questions that cannot be answered in a way that matches the reference answer without access to the gold passages, such as “What is the weather in Volcano, CA?” so performance will be lower without using gold passages. We also note that some MSMARCO questions cannot be answered using Wikipedia alone. Here, RAG can rely on parametric knowledge to generate reasonable responses.
@@ 3.3 Jeopardy Question Generation
To evaluate RAG’s generation abilities in a non-QA setting, we study open-domain question generation. Rather than use questions from standard open-domain QA tasks, which typically consist of short, simple questions, we propose the more demanding task of generating Jeopardy questions. Jeopardy is an unusual format that consists of trying to guess an entity from a fact about that entity. For example, “The World Cup” is the answer to the question “In 1986 Mexico scored as the first country to host this international sports competition twice.” As Jeopardy questions are precise, factual statements, generating Jeopardy questions conditioned on their answer entities constitutes a challenging knowledge-intensive generation task.
We use the splits from SearchQA [10], with 100K train, 14K dev, and 27K test examples. As this is a new task, we train a BART model for comparison. Following [67], we evaluate using the SQuAD-tuned Q-BLEU-1 metric [42]. Q-BLEU is a variant of BLEU with a higher weight for matching entities and has higher correlation with human judgment for question generation than standard metrics. We also perform two human evaluations, one to assess generation factuality, and one for specificity. We define factuality as whether a statement can be corroborated by trusted external sources, and specificity as high mutual dependence between the input and output [33]. We follow best practice and use pairwise comparative evaluation [34]. Evaluators are shown an answer and two generated questions, one from BART and one from RAG. They are then asked to pick one of four options—quuestion A is better, question B is better, both are good, or neither is good.
@@ 3.4 Fact Verification
FEVER [56] requires classifying whether a natural language claim is supported or refuted by Wikipedia, or whether there is not enough information to decide. The task requires retrieving evidence from Wikipedia relating to the claim and then reasoning over this evidence to classify whether the claim is true, false, or unverifiable from Wikipedia alone. FEVER is a retrieval problem coupled with an challenging entailment reasoning task. It also provides an appropriate testbed for exploring the RAG models’ ability to handle classification rather than generation. We map FEVER class labels (supports, refutes, or not enough info) to single output tokens and directly train with claim-class pairs. Crucially, unlike most other approaches to FEVER, we do not use supervision on retrieved evidence. In many real-world applications, retrieval supervision signals aren’t available, and models that do not require such supervision will be applicable to a wider range of tasks. We explore two variants: the standard 3-way classification task (supports/refutes/not enough info) and the 2-way (supports/refutes) task studied in Thorne and Vlachos [57]. In both cases we report label accuracy.
@@ 4 Results
Table 1: Open-Domain QA Test Scores. For TQA, left column uses the standard test set for Open- Domain QA, right column uses the TQA-Wiki test set. See Appendix D for further details.
@@
Open Book REALM [20] 40.4 - / - 40.7 46.8
@@
RAG-Seq. 44.5 56.8 / 68.0 45.2 52.2
Table 2: Generation and classification Test Scores. MS-MARCO SotA is [4], FEVER-3 is [68] and FEVER-2 is [57] *Uses gold context/evidence. Best model without gold access underlined.
Model Jeopardy MSMARCO FVR3 FVR2
B-1 QB-1 R-L B-1 Label Acc.
@@
RAG-Tok. 17.3 22.2 40.1 41.5 72.5 89.5
RAG-Seq. 14.7 21.4 40.8 44.2
@@ 4.1 Open-domain Question Answering
Table 2 shows results for RAG along with state-of-the-art models. On all four open-domain QA tasks, RAG sets a new state of the art (only on the T5-comparable split for TQA). RAG combines the generation flexibility of the “closed-book” (parametric only) approaches and the performance of "open-book" retrieval-based approaches. Unlike REALM and T5+SSM, RAG enjoys strong results without expensive, specialized “salient span masking” pre-training [20]. It is worth noting that RAG’s retriever is initialized using DPR’s retriever, which uses retrieval supervision on Natural Questions and TriviaQA. RAG compares favourably to the DPR QA system, which uses a BERT-based “cross-encoder” to re-rank documents, along with an extractive reader. RAG demonstrates that neither a re-ranker nor extractive reader is necessary for state-of-the-art performance.
There are several advantages to generating answers even when it is possible to extract them. Documents with clues about the answer but do not contain the answer verbatim can still contribute towards a correct answer being generated, which is not possible with standard extractive approaches, leading to more effective marginalization over documents. Furthermore, RAG can generate correct answers even when the correct answer is not in any retrieved document, achieving 11.8% accuracy in such cases for NQ, where an extractive model would score 0%.
@@ 4.2 Abstractive Question Answering
As shown in Table 2, RAG-Sequence outperforms BART on Open MS-MARCO NLG by 2.6 Bleu points and 2.6 Rouge-L points. RAG approaches state-of-the-art model performance, which is impressive given that (i) those models access gold passages with specific information required to generate the reference answer , (ii) many questions are unanswerable without the gold passages, and (iii) not all questions are answerable from Wikipedia alone. Table 3 shows some generated answers from our models. Qualitatively, we find that RAG models hallucinate less and generate factually correct text more often than BART. Later, we also show that RAG generations are more diverse than BART generations (see §4.5).
@@ 4.3 Jeopardy Question Generation
Table 2 shows that RAG-Token performs better than RAG-Sequence on Jeopardy question generation, with both models outperforming BART on Q-BLEU-1. 5 shows human evaluation results, over 452 pairs of generations from BART and RAG-Token. Evaluators indicated that BART was more factual than RAG in only 7.1% of cases, while RAG was more factual in 42.7% of cases, and both RAG and BART were factual in a further 17% of cases, clearly demonstrating the effectiveness of RAG on the task over a state-of-the-art generation model. Evaluators also find RAG generations to be more specific by a large margin. Table 3 shows typical generations from each model.
Jeopardy questions often contain two separate pieces of information, and RAG-Token may perform best because it can generate responses that combine content from several documents. Figure 2 shows an example. When generating “Sun”, the posterior is high for document 2 which mentions “The Sun Also Rises”. Similarly, document 1 dominates the posterior when “A Farewell to Arms” is generated. Intriguingly, after the first token of each book is generated, the document posterior flattens. This observation suggests that the generator can complete the titles without depending on specific documents. In other words, the model’s parametric knowledge is sufficient to complete the titles. We find evidence for this hypothesis by feeding the BART-only baseline with the partial decoding "The Sun. BART completes the generation "The Sun Also Rises" is a novel by this author of "The Sun Also Rises" indicating the title "The Sun Also Rises" is stored in BART’s parameters. Similarly, BART will complete the partial decoding "The Sun Also Rises" is a novel by this author of "A with "The Sun Also Rises" is a novel by this author of "A Farewell to Arms". This example shows how parametric and non-parametric memories work together—the non-parametric component helps to guide the generation, drawing out specific knowledge stored in the parametric memory.
Figure 2: RAG-Token document posterior $p(z_{i}|x,y_{i},y_{-i})$ for each generated token for input “Hemingway" for Jeopardy generation with 5 retrieved documents. The posterior for document 1 is high when generating “A Farewell to Arms" and for document 2 when generating “The Sun Also Rises".
Table 3: Examples from generation tasks. RAG models generate more specific and factually accurate responses. ‘?’ indicates factually incorrect responses, * indicates partially correct responses.
@@
Factuality Specificity
BART better 7.1% 16.8%
RAG better 42.7% 37.4%
Both good 11.7% 11.8%
Both poor 17.7% 6.9%
No majority 20.8% 20.1%
Table 5: Ratio of distinct to total tri-grams for generation tasks.
MSMARCO Jeopardy QGen
Gold 89.6% 90.0%
BART 70.7% 32.4%
RAG-Token 77.8% 46.8%
RAG-Seq. 83.5% 53.8%
@@ 4.4 Fact Verification
Table 2 shows our results on FEVER. For 3-way classification, RAG scores are within 4.3% of state-of-the-art models, which are complex pipeline systems with domain-specific architectures and substantial engineering, trained using intermediate retrieval supervision, which RAG does not require. For 2-way classification, we compare against Thorne and Vlachos [57], who train RoBERTa [35] to classify the claim as true or false given the gold evidence sentence. RAG achieves an accuracy within 2.7% of this model, despite being supplied with only the claim and retrieving its own evidence. We also analyze whether documents retrieved by RAG correspond to documents annotated as gold evidence in FEVER. We calculate the overlap in article titles between the top $k$ documents retrieved by RAG and gold evidence annotations. We find that the top retrieved document is from a gold article in 71% of cases, and a gold article is present in the top 10 retrieved articles in 90% of cases.
4.5 Additional Results
Generation Diversity
Section 4.3 shows that RAG models are more factual and specific than BART for Jeopardy question generation. Following recent work on diversity-promoting decoding [33, 59, 39], we also investigate generation diversity by calculating the ratio of distinct ngrams to total ngrams generated by different models. Table 5 shows that RAG-Sequence’s generations are more diverse than RAG-Token’s, and both are significantly more diverse than BART without needing any diversity-promoting decoding.
Table 6: Ablations on the dev set. As FEVER is a classification task, both RAG models are equivalent.
@@ 4.5 Additional Results
RAG-Sequence 44.0 55.8 44.9 53.4 15.3 21.5 57.2 47.5
Retrieval Ablations
A key feature of RAG is learning to retrieve relevant information for the task. To assess the effectiveness of the retrieval mechanism, we run ablations where we freeze the retriever during training. As shown in Table 6, learned retrieval improves results for all tasks.
@@
We compare RAG’s dense retriever to a word overlap-based BM25 retriever [53]. Here, we replace RAG’s retriever with a fixed BM25 system, and use BM25 retrieval scores as logits when calculating $p(z|x)$. Table 6 shows the results. For FEVER, BM25 performs best, perhaps since FEVER claims are heavily entity-centric and thus well-suited for word overlap-based retrieval. Differentiable retrieval improves results on all other tasks, especially for Open-Domain QA, where it is crucial.
Index hot-swapping
@@
An advantage of non-parametric memory models like RAG is that knowledge can be easily updated at test time. Parametric-only models like T5 or BART need further training to update their behavior as the world changes. To demonstrate, we build an index using the DrQA [5] Wikipedia dump from December 2016 and compare outputs from RAG using this index to the newer index from our main results (December 2018). We prepare a list of 82 world leaders who had changed between these dates and use a template “Who is {position}?” (e.g. “Who is the President of Peru?”) to query our NQ RAG model with each index. RAG answers 70% correctly using the 2016 index for 2016 world leaders and 68% using the 2018 index for 2018 world leaders. Accuracy with mismatched indices is low (12% with the 2018 index and 2016 leaders, 4% with the 2016 index and 2018 leaders). This shows we can update RAG’s world knowledge by simply replacing its non-parametric memory.
Effect of Retrieving more documents
Models are trained with either 5 or 10 retrieved latent documents, and we do not observe significant differences in performance between them. We have the flexibility to adjust the number of retrieved documents at test time, which can affect performance and runtime. Figure 3 (left) shows that retrieving more documents at test time monotonically improves Open-domain QA results for RAG-Sequence, but performance peaks for RAG-Token at 10 retrieved documents. Figure 3 (right) shows that retrieving more documents leads to higher Rouge-L for RAG-Token at the expense of Bleu-1, but the effect is less pronounced for RAG-Sequence.
Figure 3: Left: NQ performance as more documents are retrieved. Center: Retrieval recall performance in NQ. Right: MS-MARCO Bleu-1 and Rouge-L as more documents are retrieved.